# A Regime-Aware Fusion Framework for Time Series Classification

### v1

@@ 2 Related Work
Using raw sequences with generic distances (e.g., Euclidean) is brittle under noise, scaling, and time warping [Lin2007SAX]. A central design choice in TSC is therefore the representation: a transformation that exposes structure useful for discrimination. Large empirical studies—most notably the “Great Time Series Classification Bake Off” [Bagnall2017BakeOff] and its recent follow-up [Middlehurst2021BakeoffRedux]—show that no single representation dominates across datasets, motivating families of complementary transformations.
The TSC literature offers a rich taxonomy of representations. Shapelets identify discriminative local subsequences via information gain or distance-based scoring [Hills2014]. Catch22 provides a compact set of 22 canonical time series features selected from over 7000 candidates for broad domain coverage [Lubba2019Catch22]. Continuous Wavelet Transform (CWT) decomposes signals into time-frequency representations, useful for non-stationary patterns. MultiROCKET extends ROCKET with additional pooling statistics and multi-resolution kernels [Tan2022MultiROCKET]. Dictionary-based methods (e.g., BOSS, cBOSS) combine symbolic discretisation with bag-of-patterns classifiers [Schafer2015BOSS]. Deep learning approaches—CNNs, ResNets, InceptionTime, and Transformers—learn end-to-end hierarchical features but typically require more data and compute [Fawaz2019Review, Ismail2020InceptionTime].
@@
SAX maps -normalised segment means (PAA) to symbols via Gaussian breakpoints, enabling lower bounds and motif discovery. Its behaviour depends on the windowing and alphabet parameters. SFA instead truncates local DFTs and discretises each coefficient via Multiple Coefficient Binning, capturing global/spectral regularities and shift tolerance, often complementary to time-domain cues [Schafer2012SFA]. ROCKET replaces learned CNN filters with thousands of random kernels of varied lengths/dilations and summarises each response by max and PPV; a linear/ridge classifier then operates on these features [Dempster2020ROCKET]. The result is near-state-of-the-art accuracy with excellent speed and scalability. MiniROCKET [Dempster2021MiniROCKET] and MultiROCKET [Tan2022MultiROCKET] refine this recipe, but keep the same principle: diverse random convolutions + cheap pooling.
Because dataset characteristics vary, a long-standing question is how to identify the most discriminative representation [Abanda2009]. Shapelet methods target local, discriminative motifs and perform well on some domains (e.g., ECG/outline datasets), but broad evaluations show no single approach dominates across datasets [Middlehurst2021BakeoffRedux, Bagnall2017BakeOff]. Prior work has compared transformations via: (i) distance fidelity (e.g., TLB for lower-bounded DTW surrogates) [Wang2013TLB]; and (ii) global statistical criteria (e.g., information gain, -tests, Kruskal–Wallis) [Hills2014]. Empirical studies also emphasise that performance hinges more on features than on the downstream classifier [Nakano2014].
Most comparisons operate at the dataset-level. High TLB or superior average accuracy does not explain when and why a representation helps at the instance level, nor how to adapt across heterogeneous data within a dataset [Abanda2009, Serra2015Similarity]. Moreover, non-adaptive ensembles aggregate evidence but rarely selectively prioritise the most informative representation per sample in a way that is both effective and interpretable.
@@ 3.1 Representations and Fusion
We use Sax (symbolic time-domain), Sfa (symbolic frequency-domain), and Rocket (random convolutional kernels). This choice is justified by both theoretical and empirical evidence. Primarily, our choice spans complementary domains and resolutions. SFA operates in the frequency domain via truncated DFT and coefficient binning, exposing spectral periodicities and shift-invariant global structure. SAX and ROCKET both operate in the time domain, but capture fundamentally different aspects of temporal structure: SAX provides coarse, noise-resistant symbolic summaries of segment-level trends (e.g., “rising then flat”), whereas ROCKET extracts fine-grained, high-resolution local shapes and transients via thousands of random convolutional kernels (e.g., specific spike patterns, edge responses). This multi-resolution temporal coverage is a deliberate design choice validated by state-of-the-art ensemble methods—HIVE-COTE [Lines2018HIVECOTE] and Mr-Hydra [Dempster2023Hydra] explicitly combine multiple time-domain representations like shapelet/dictionary-based methods (akin to SAX’s symbolic coarseness) with convolutional features (akin to ROCKET’s fine-grained kernels) to achieve top performance. Empirically, the correlation matrix of accuracy ranks from benchmark studies [Bagnall2017BakeOff, Middlehurst2021BakeoffRedux] confirms that SAX-like and ROCKET-like classifiers are often negatively correlated or weakly correlated, indicating they excel on different subsets of datasets.
@@ 4.1 Hyperparameter Search and Fairness (Full Grid)
To ensure comparability, we used a full grid over small, literature-backed [Middlehurst2021BakeoffRedux] ranges and applied the same head capacity and training schedule to all models. SAX: word , frame , alphabet ; SFA: word , window , alphabet ; ROCKET: , seed ; Head (all models): embed/hidden , dropout ; Training: LR , batch , max epochs with patience , folds, seed . These ranges span coarsefine granularity in both time and frequency while matching prior studies [Bagnall2017BakeOff, Middlehurst2021BakeoffRedux] and keeping cost tractable. Pilots showed broader ranges yielded pp median gain but – runtime. Fixing the head and schedule prevents capacity/training time from confounding representation comparisons. Search depth: F3 ( configs/dataset), F2_SR/F2_SFR (), SAX/SFA solo (), ROCKET solo (). Using a full grid ensures each model family receives the same search depth; training over HPC infrastructure made this feasible on all 113 datasets.
@@ 4.2 Meta-Features and Regimes
We compute a range of 13 low-correlated dataset meta-features categorised into two groups: Global complexity features: series length, turning points and variance, spectral entropy and its variance, KL divergence of the power spectrum, permutation entropy, autocorrelation lag-1 and kurtosis. Class separability features: DTW separability time and frequency domain, Kruskal power spectral density of classes and imbalance index. This grouping reflects two complementary logics. Global features describe the intrinsic properties of the time series themselves (e.g., entropy, length), which are useful for identifying broad structural similarities across datasets. Class-based metrics, in contrast, explicitly exploit label information to measure separability (e.g., DTW class distances). Including both ensures that clustering is informed by how datasets look in general, as well as how hard they are to separate in practice. Feature choice was guided by prior work and practical utility. DTW-based separability follows Wang et al. [Wang2013TLB], where lower-bound distances (LBKeogh) were shown as strong quality indicators. Statistical measures such as Kruskal PSD align with Hills et al. [Hills2014], which emphasised distributional/statistical descriptors. Dataset-level factors like imbalance are well known to affect classification difficulty. Variance-based counterparts (e.g., turning_points_var, kurtosis_var, spectral_entropy_var) were included to capture intra-dataset volatility. While averages capture central tendencies, variance reflects whether a property is consistent across all series or dominated by a few irregular ones. Capturing both aspects provides a richer description of the dataset structure. Taken together, the final feature set spans global, class-based, and variance-sensitive perspectives, providing a balanced and logically grounded foundation for clustering. Detailed mathematical definitions and formulas for all meta-features are provided in Appendix C, Tables 9 and 10. Hierarchical clustering yields six regimes: C1 HighImb(high imbalance signals), C2 LongFSTCx (long, frequency-separable time-complex signals), C3 SmoothSep(smooth and separable signals), C4 HighFlCx (highly fluctuating complex signals), C5 HighCompOut(high complexity outlier rich signals), C6 ShortBase(short baseline signals).
@@ Acknowledgments
We thank the University of Bristol Advanced Computing Research Centre for providing HPC resources that made the extensive hyperparameter grid searches feasible across all 113 datasets. We also acknowledge the UCR Time Series Classification Archive [Dau2019UCR] for providing the benchmark datasets used in this study.
@@ C. Meta-Features
| Meta-Feature | Formula | Level | Indicator (higher ) |
| spectral_entropy_glb | Global (per-series mean) | More noise / complexity | |
| spectral_entropy_var_glb | Global | Greater diversity of complexity | |
| turning_points_glb | Global (per-series mean) | More volatile / oscillatory | |
| turning_points_var_glb | Global | Greater volatility heterogeneity | |
| kurtosis_glb | Global (per-series mean) | Heavier tails / more spikes | |
| autocorr_lag1_glb | Global (per-series mean) | Stronger short-term memory | |
| permutation_entropy_glb | (order , delay ) | Global (per-series mean) | More irregular / unpredictable |
| ts_length_glb | Global | Longer temporal context | |
| kl_psd_glb | Global | Greater spectral diversity | |
| dtw_separability_time_cls | Class-based (subsampled) | Better time-domain separability | |
| dtw_separability_freq_cls | Same as above but on PSD sequences | Class-based (subsampled) | Better frequency-domain separability |
| kruskal_psd_cls | Class-based | Stronger class spectral differences | |
| imbalance_index_cls | (default accuracy) | Class-based | Stronger class imbalance |
Notes: All per-series features computed on z-normalized series, then averaged across samples. Welch PSD uses ; logs are natural. For DTW-based features, expectations are computed on a dynamically selected, approximately balanced subsample per class. The per-dataset budget is
with and the series length, for efficiency and comparability across datasets.
| Feature Group | Features and Justification |
| Shape / Statistics |
turning_points, turning_points_var: capture local oscillations and variability in signal shape.
kurtosis: tailedness/peakedness; detect outliers and sporadic spikes. autocorr_lag1: short-term temporal dependence. |
| Spectral / Entropy |
spectral_entropy, spectral_entropy_var: frequency complexity and heterogeneity.
permutation_entropy: temporal unpredictability (order-pattern randomness). kruskal_psd, global_kl_psd: class-wise PSD differences and overall spectral diversity. |
| Separability Measures | dtw_separability_time, dtw_separability_freq: DTW separability in time/frequency domains. |
| Dataset Properties |
ts_length: average series length (temporal context).
imbalance: class distribution skewness (default accuracy). |

### v2

@@ 2 Related Work
Using raw sequences with generic distances (e.g., Euclidean) is brittle under noise, scaling, and time warping [11]. A central design choice in TSC is therefore the representation: a transformation that exposes structure useful for discrimination. Large empirical studies—most notably the “Great Time Series Classification Bake Off” [3] and its recent follow-up [14]—show that no single representation dominates across datasets, motivating families of complementary transformations.
The TSC literature offers a rich taxonomy of representations. Shapelets identify discriminative local subsequences via information gain or distance-based scoring [9]. Catch22 provides a compact set of 22 canonical time series features selected from over 7000 candidates for broad domain coverage [13]. Continuous Wavelet Transform (CWT) decomposes signals into time-frequency representations, useful for non-stationary patterns. MultiROCKET extends ROCKET with additional pooling statistics and multi-resolution kernels [19]. Dictionary-based methods (e.g., BOSS, cBOSS) combine symbolic discretisation with bag-of-patterns classifiers [17]. Deep learning approaches—CNNs, ResNets, InceptionTime, and Transformers—learn end-to-end hierarchical features but typically require more data and compute [8, 10].
@@
SAX maps -normalised segment means (PAA) to symbols via Gaussian breakpoints, enabling lower bounds and motif discovery. Its behaviour depends on the windowing and alphabet parameters. SFA instead truncates local DFTs and discretises each coefficient via Multiple Coefficient Binning, capturing global/spectral regularities and shift tolerance, often complementary to time-domain cues [16]. ROCKET replaces learned CNN filters with thousands of random kernels of varied lengths/dilations and summarises each response by max and PPV; a linear/ridge classifier then operates on these features [5]. The result is near-state-of-the-art accuracy with excellent speed and scalability. MiniROCKET [6] and MultiROCKET [19] refine this recipe, but keep the same principle: diverse random convolutions + cheap pooling.
Because dataset characteristics vary, a long-standing question is how to identify the most discriminative representation [1]. Shapelet methods target local, discriminative motifs and perform well on some domains (e.g., ECG/outline datasets), but broad evaluations show no single approach dominates across datasets [14, 3]. Prior work has compared transformations via: (i) distance fidelity (e.g., TLB for lower-bounded DTW surrogates) [20]; and (ii) global statistical criteria (e.g., information gain, -tests, Kruskal–Wallis) [9]. Empirical studies also emphasise that performance hinges more on features than on the downstream classifier [15].
Most comparisons operate at the dataset-level. High TLB or superior average accuracy does not explain when and why a representation helps at the instance level, nor how to adapt across heterogeneous data within a dataset [1, 18]. Prior work has demonstrated the benefits of representation cooperation through ensemble methods that combine complementary symbolic representations, such as Co-eye [2], which uses multiple hyper-parameterised symbolic representations (PAA and Fourier approximations) with soft dynamic voting. However, non-adaptive ensembles aggregate evidence but rarely selectively prioritise the most informative representation per sample in a way that is both effective and interpretable.
@@ 3.1 Representations and Fusion
We use Sax (symbolic time-domain), Sfa (symbolic frequency-domain), and Rocket (random convolutional kernels). This choice is justified by both theoretical and empirical evidence. Primarily, our choice spans complementary domains and resolutions. SFA operates in the frequency domain via truncated DFT and coefficient binning, exposing spectral periodicities and shift-invariant global structure. SAX and ROCKET both operate in the time domain, but capture fundamentally different aspects of temporal structure: SAX provides coarse, noise-resistant symbolic summaries of segment-level trends (e.g., “rising then flat”), whereas ROCKET extracts fine-grained, high-resolution local shapes and transients via thousands of random convolutional kernels (e.g., specific spike patterns, edge responses). This multi-resolution temporal coverage is a deliberate design choice validated by state-of-the-art ensemble methods—HIVE-COTE [12] and Mr-Hydra [7] explicitly combine multiple time-domain representations like shapelet/dictionary-based methods (akin to SAX’s symbolic coarseness) with convolutional features (akin to ROCKET’s fine-grained kernels) to achieve top performance. Empirically, the correlation matrix of accuracy ranks from benchmark studies [3, 14] confirms that SAX-like and ROCKET-like classifiers are often negatively correlated or weakly correlated, indicating they excel on different subsets of datasets.
@@ 4.1 Hyperparameter Search and Fairness (Full Grid)
To ensure comparability, we used a full grid over small, literature-backed [14] ranges and applied the same head capacity and training schedule to all models. SAX: word , frame , alphabet ; SFA: word , window , alphabet ; ROCKET: , seed ; Head (all models): embed/hidden , dropout ; Training: LR , batch , max epochs with patience , folds, seed . These ranges span coarsefine granularity in both time and frequency while matching prior studies [3, 14] and keeping cost tractable. Pilots showed broader ranges yielded pp median gain but – runtime. Fixing the head and schedule prevents capacity/training time from confounding representation comparisons. Search depth: F3 ( configs/dataset), F2_SR/F2_SFR (), SAX/SFA solo (), ROCKET solo (). Using a full grid ensures each model family receives the same search depth; training over HPC infrastructure made this feasible on all 113 datasets.
@@ 4.2 Meta-Features and Regimes
We compute a range of 13 low-correlated dataset meta-features categorised into two groups: Global complexity features: series length, turning points and variance, spectral entropy and its variance, KL divergence of the power spectrum, permutation entropy, autocorrelation lag-1 and kurtosis. Class separability features: DTW separability time and frequency domain, Kruskal power spectral density of classes and imbalance index. This grouping reflects two complementary logics. Global features describe the intrinsic properties of the time series themselves (e.g., entropy, length), which are useful for identifying broad structural similarities across datasets. Class-based metrics, in contrast, explicitly exploit label information to measure separability (e.g., DTW class distances). Including both ensures that clustering is informed by how datasets look in general, as well as how hard they are to separate in practice. Feature choice was guided by prior work and practical utility. DTW-based separability follows Wang et al. [20], where lower-bound distances (LBKeogh) were shown as strong quality indicators. Statistical measures such as Kruskal PSD align with Hills et al. [9], which emphasised distributional/statistical descriptors. Dataset-level factors like imbalance are well known to affect classification difficulty. Variance-based counterparts (e.g., turning_points_var, kurtosis_var, spectral_entropy_var) were included to capture intra-dataset volatility. While averages capture central tendencies, variance reflects whether a property is consistent across all series or dominated by a few irregular ones. Capturing both aspects provides a richer description of the dataset structure. Taken together, the final feature set spans global, class-based, and variance-sensitive perspectives, providing a balanced and logically grounded foundation for clustering. Detailed mathematical definitions and formulas for all meta-features are provided in Appendix C, Tables 9 and 10. Hierarchical clustering yields six regimes: C1 HighImb(high imbalance signals), C2 LongFSTCx (long, frequency-separable time-complex signals), C3 SmoothSep(smooth and separable signals), C4 HighFlCx (highly fluctuating complex signals), C5 HighCompOut(high complexity outlier rich signals), C6 ShortBase(short baseline signals).
@@ Acknowledgments
We thank the University of Bristol Advanced Computing Research Centre for providing HPC resources that made the extensive hyperparameter grid searches feasible across all 113 datasets. We also acknowledge the UCR Time Series Classification Archive [4] for providing the benchmark datasets used in this study.
@@ C. Meta-Features
| Meta-Feature | Formula | Level | Higher values indicate |
| spectral_entropy∗ | Global* | More noise-like / complex spectra | |
| spectral_entropy_var | Global | Greater heterogeneity in spectral complexity | |
| turning_points∗ | Global* | More oscillatory local structure | |
| turning_points_var | Global | Heterogeneous local volatility | |
| kurtosis∗ | Global* | Heavier tails / spike-like behavior | |
| autocorr_lag1∗ | Global* | Stronger short-term temporal dependence | |
| permutation_entropy∗ | Global* | Greater ordinal irregularity | |
| ts_length | Global | Longer temporal context | |
| kl_psd∗ | Global* | Higher global spectral diversity | |
| dtw_separability_time | Class-based | Easier time-domain class separation | |
| dtw_separability_freq | Same ratio applied to PSD sequences | Class-based | Easier frequency-domain separation |
| kruskal_psd | on PSD energy by class | Class-based | Stronger class spectral differences |
| imbalance_index | Class-based | Stronger class imbalance |
Notes. Features marked with ∗ are computed per time series after z-normalization and then averaged across samples to obtain a dataset-level summary. Variance features are computed directly across per-series values and require no additional aggregation. Time series length is constant within each dataset and is therefore treated as a dataset-level attribute. Power spectral density (PSD) estimates use Welch’s method with and natural logarithms. Class-based DTW separability estimates and via averages over approximately balanced, dynamically subsampled class pairs, reducing dominance of large classes and ensuring comparability across datasets.
| Feature Group | Included features and rationale |
| Shape / Statistics | turning_points, turning_points_var: capture local oscillatory structure and heterogeneity; kurtosis: identifies heavy-tailed or spike-like behavior; autocorr_lag1: measures short-term temporal dependence. |
| Spectral / Entropy | spectral_entropy, spectral_entropy_var: quantify frequency-domain complexity and diversity; permutation_entropy: measures ordinal unpredictability; kruskal_psd, global_kl_psd: capture class-wise spectral separation and global spectral diversity. |
| Separability Measures | dtw_separability_time, dtw_separability_freq: quantify class separability in the time and frequency domains using DTW-based distance ratios. |
| Dataset Properties | ts_length: reflects available temporal context; imbalance: captures class distribution skewness (default accuracy baseline). |